OpenStudy (jack1):

ok i have a question please: how do i graph (online) a variable equation: ie - if x > 100, y = 200/x^2, but if x < 100, y = 150/ x^2 specifically my 2 equations are: y = (1.5417x10^-6) x, for x = 0 to 6371000 y = 3.98678x10^14/x^2, for x = 637100 to 10000000

ganeshie8 (ganeshie8):

geogebra its easy to plot piecewise functions

OpenStudy (jack1):

nice one, cheers man !

ganeshie8 (ganeshie8):

u have geogebra ?

OpenStudy (jack1):

nope, but about to download

ganeshie8 (ganeshie8):

cool , download, its free + takes oly few seconds to install

OpenStudy (jack1):

did u get the chrome plugin or did u go for the windows/ apple install option?

ganeshie8 (ganeshie8):

install..

ganeshie8 (ganeshie8):

http://www.geogebra.org/download/?os=win

ganeshie8 (ganeshie8):

thats for windows

ganeshie8 (ganeshie8):

for piecewise function :- y = 1 ; x > 10 = 0 ; x < 10 enter below :- f(x) = If[x > 10, 1,if[x < 10, 0]]

OpenStudy (jack1):

hmm, plotted... is there a zoom button dude or a : display range...?

ganeshie8 (ganeshie8):

use ctr+, ctrl- like u do in chrome

OpenStudy (jack1):

ahh, brilliant, cheers man this is tops hey

ganeshie8 (ganeshie8):

you may plot it online in desmos calculator aswell uve used it before right ?

ganeshie8 (ganeshie8):

https://www.desmos.com/calculator instructions for plotting piecewise funcs http://www.youtube.com/watch?v=L5P8l7VDncc&feature=plcp

OpenStudy (jack1):

no mate, never hey, have bookmarked this q so i can use all the tools now and remember the piecewise function input

OpenStudy (jack1):

lol

ganeshie8 (ganeshie8):

lol, desmos is an online graphing tool... its really easy... try entering below : y = {x<10:0, x>10:1} https://www.desmos.com/calculator

OpenStudy (jack1):

no go dude... i think i have the correct parameters in, but it's giving me no results

OpenStudy (jack1):

here's what i plugged into desmos $f\left(x\right)=\space \left\{x<6371000:1.5417\cdot 10^{\left(-6\right)}x,\space x>6371000:3.98678\cdot 10^{\left(14\right)}/x^2\right\}$

OpenStudy (jack1):

damn, that doesn't look right

ganeshie8 (ganeshie8):

for small values of x, output is very low. pfa screenshot... to see some y value, u need to go right atleast upto x = 6371000

OpenStudy (jack1):

thanks man, yeah i managed to zoom in the y axis on the geogebra program, cheers for the follow up tho man